OpenStudy (purexyz):

Write an equation of a line in slope intercept form that is perpendicular to y= -4x-2 passes through (-16,-11) .

OpenStudy (anonymous):

If we want 2 line they are perpendicular they slope are this : m2 = -1/m1 For example : Find an equation perpendicular to y = 2x +8 We just find that slope line : That slope is -1/2 So we have : y=-1/2+8 Just this . Got it ?

OpenStudy (purexyz):

Can you explain a little bit more please?

OpenStudy (anonymous):

And if we want that line through a point help from this : Y-y2=m(X2-x) x2 and y2 are the numbers we have in the Coordinate s line . And x and y are in the form . So in the example if we want that line through (-5 , 4) have : y=-1/2+8 is our equation use from : Y-y2=m(X2-x) have : y-4=-1/2(x-(-5))=>y-4=-1/2(x+5) and write it in standard form it s y=mx+d means : y=-1/2x-5/2-4 y=-1/2-13/2

OpenStudy (anonymous):

Of course ! Look at this example . If we want a line perpendicular to another line :  (line d )

OpenStudy (anonymous):

have : 

OpenStudy (anonymous):

that equation certainly have a slope that slope is Symmetric inverse . Means if the slope for d is 2 for the black one is -1/2 . Got it ?

OpenStudy (purexyz):

yeah thanks

OpenStudy (anonymous):

And we want the black line for example passes through (2 , 1) use this : y2-y=m(x2-x) y2=1 x2=2 m=-1/2 1-y=-1/2(2-x) ( I think had a fall in the first example there ) 1-y=-1+1/2x -y=-1-1+1/2x y=+2-1/2x And this is our end equation .

OpenStudy (anonymous):

For your question you should find slope(m) first and use from : y2-y=m(x2-x) X2=-16 Y2=-11

OpenStudy (anonymous):

@E.ali you have you equation wrong it y-y1=m(x-x1) y-(-16)=1/4(x-(-11)) Distribute y+16=1/4x+2.75 Subtract 16 y=1/4x-13.25